5. Rationalise the denominators of the following :
(i) 1 (ii) 7 1
7 - 6
(iii) 1 (iv) 1
5 + 2 7 - 2



Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) 1 7 Multiplying the numerator and denominator by 7 we have 1 = 1 # 7 = 7 7 7 7 7 (ii) 1 7 - 6 Multiplying the numerator and denominator by ( 7 + 6 ) we have 1 = 7 + 6 7 - 6 ( 7− 6)( 7+ 6) = 7 + 6 = 7 + 6 7 − 6 (iii) 1 5 + 2 Multiplying the numerator and denominator by ( 5 - 2 ) we have 1 = 5 − 2 5 + 2 ( 5+ 2)( 5− 2) = 5− 2= 5− 2 5−23 (iv) 1 7 - 2 Multiplying the numerator and denominator by ( 7+2) we have 1 = 7 + 2 7 - 2 ( 7 −2)( 7 +2) Support Us = 77−+42 = 73+ 2